the perimeter of a rectangle is 96 and the length is 6 meters longer than twice the width. what are the length and width
Answer by josgarithmetic(39799)

Perimeter & rectangle dimensions questions like this are too typical.

L for length
w for width

L=2w+6 and 2w+2L=96. These are two equations in two unknowns, and easily solved.

The perimeter equation in this example simplifies to w+L=48.
Substitute: w+(2w+6)=48
3w=48-6
3w=42
w=(6*7)/3
w=14--------------one of the answers.

L=2*14+6 = ?